OpenStudy (anonymous):

Solve 2x^2 = −10x by factoring possible answers: x = 2, x = −5 x = 2, x = 5 x = 0, x = 5 x = 0, x = −5

OpenStudy (anonymous):

First get the -10x on the left by adding it to both sides

OpenStudy (anonymous):

ok. then?

OpenStudy (anonymous):

You have 2x^2+10x=0 right?

OpenStudy (anonymous):

What can you factor out of 2x^2+10x?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

Yep, so you get 2x(x+5)=0

OpenStudy (anonymous):

So 2x=0 and x+5=0 solve for x in both of those

OpenStudy (anonymous):

ok. so x=0 and x=-5?

OpenStudy (anonymous):

Yep thats correct

OpenStudy (anonymous):

ok thanks:)

OpenStudy (anonymous):

You're welcome